OpenStudy (anonymous):

what is the solution to the rational equation, x/2x-1 - 1/4 = 2/2x-1

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

i found the least common factor and tried that but my answer didnt come out right

OpenStudy (anonymous):

hey Emily wat do u need help on

OpenStudy (skullpatrol):

$$\huge \frac{x}{2x-1} - \frac{1}{4} = \frac{2}{2x-1}$$

OpenStudy (skullpatrol):

Is that^ the question?

OpenStudy (anonymous):

yes! that ittt

OpenStudy (anonymous):

wat r the answer choices

OpenStudy (anonymous):

the answer choices are x=2 x=7/2 x=3 x=7/6

OpenStudy (anonymous):

Unbock me and I will help u

OpenStudy (anonymous):

I want 2 help u and i can't if u don't unblock me

OpenStudy (anonymous):

i did unblock youu

OpenStudy (anonymous):

^^thank youuuuu

OpenStudy (anonymous):

no problem

OpenStudy (luigi0210):

@kathy0514 no giving away answers

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

Let the student learn please...

OpenStudy (skullpatrol):

@emilyporazinski are you still there?